# Discovering Concept Directions from Diffusion-based Counterfactuals via Latent Clustering

### v1

@@ Abstract
Concept-based explanations have emerged as an effective approach within Explainable Artificial Intelligence, enabling interpretable insights by aligning model decisions with human-understandable concepts. However, existing methods rely on computationally intensive procedures and struggle to efficiently capture complex, semantic concepts. Recently, the Concept Discovery through Latent Diffusion-based Counterfactual Trajectories (CDCT) framework, introduced by Varshney et al. (2025), attempts to identify concepts via dimension-wise traversal of the latent space of a Variational Autoencoder trained on counterfactual trajectories. Extending the CDCT framework, this work introduces Concept Directions via Latent Clustering (CDLC), which extracts global, class-specific concept directions by clustering latent difference vectors derived from factual and diffusion-generated counterfactual image pairs. CDLC substantially reduces computational complexity by eliminating the exhaustive latent dimension traversal required in CDCT and enables the extraction of multidimensional semantic concepts encoded across the latent dimensions. This approach is validated on a real-world skin lesion dataset, demonstrating that the extracted concept directions align with clinically recognized dermoscopic features and, in some cases, reveal dataset-specific biases or unknown biomarkers. These results highlight that CDLC is interpretable, scalable, and applicable across high-stakes domains and diverse data modalities.
@@ 1 Introduction
In high-stakes applications, such as medical diagnosis, financial risk assessment, and autonomous driving, understanding the rationale behind a neural network’s decision is often as important as the decision itself. Explainable Artificial Intelligence (XAI) [1, 2] has therefore emerged as a critical research area, aiming to bridge the gap between high-performing black-box models and human interpretability. Among the various XAI paradigms, concept-based explanations [3, 4] have gained particular attention due to their ability to express model behavior in terms of high-level, semantically meaningful concepts, rather than low‑level feature weights or pixel-based saliency maps [5, 6]. By aligning explanations with concepts recognized by domain experts, these methods facilitate trust [7, 8], debugging [9], and regulatory compliance [10, 11].
@@
Although concept-based explainability has been widely explored using convolutional [3, 4] and GAN-based architectures [12, 13], its application within diffusion-based generative models remains relatively underexplored. Recent works, such as Concept-Guided Latent Diffusion Counterfactual Explanations (CoLa-DCE) [14], have demonstrated the ability to produce spatially constrained, concept-conditioned counterfactuals for an arbitrary classifier. However, this approach only utilizes concepts to guide counterfactual explanations and does not extract global concept representations that are generalizable across examples.
@@
To discover global concepts that generalize across multiple samples, Varshney et al. [15] proposed the Concept Discovery through Latent Diffusion-based Counterfactual Trajectories (CDCT) framework. CDCT generates a classifier-guided counterfactual image trajectory dataset using a latent diffusion model [16] and subsequently trains a Variational Autoencoder (VAE) [17] on this trajectory dataset to disentangle classifier-relevant concepts. While CDCT represents a significant advancement in leveraging diffusion-based counterfactuals for concept discovery, it relies on a dimension-wise traversal strategy, wherein each latent variable is modified independently to detect relevant concepts. This exhaustive search procedure is computationally expensive, particularly in high-dimensional latent spaces, and inherently overlooks semantic concepts that arise from interactions among multiple latent dimensions. Consequently, CDCT often identifies only simple concepts, restricting its capacity to uncover high-level semantic directions.
@@
To address these limitations, this paper proposes a novel framework Concept Directions via Latent Clustering (CDLC), that extracts multi-dimensional concepts by clustering latent difference vectors computed from factual–counterfactual image pairs. These difference vectors, derived from VAE encodings of factual images and their corresponding diffusion-based counterfactuals, capture the classifier-induced transformation in latent space. In contrast to CDCT, which modifies individual latent dimensions, CDLC leverages directional clustering to reveal coordinated latent changes that correspond to semantically meaningful concepts. This approach not only reduces computational complexity by eliminating exhaustive per-dimension search but also enables the discovery of classifier-relevant concept directions that emerge from interactions among multiple latent dimensions, effectively overcoming a limitation of the original CDCT formulation.
@@
    A novel framework, Concept Directions via Latent Clustering (CDLC), is introduced as an extension of CDCT [15], significantly reducing computational complexity compared to dimension-wise search.
@@ 2 Related Work
Explainable Artificial Intelligence (XAI) aims to make machine learning models more transparent by providing human-understandable insights into their decision-making processes [1, 2]. Among various XAI approaches, concept-based explanations have gained increasing attention due to their ability to link internal model representations to semantically meaningful, human-interpretable concepts [3, 4, 18]. In supervised settings, concept-based methods rely on expert-provided annotations, which are costly and time-consuming to collect [3, 19]. For instance, Testing with Concept Activation Vectors (TCAV) [3] requires example sets for each concept and learns linear concept activation vectors to quantify a model’s sensitivity to those concepts. While effective, this reliance on curated concept examples limits the scalability and generalization of such methods. To overcome these limitations, unsupervised concept-based methods have been introduced [4, 20]. A prominent example is Automatic Concept Explanations (ACE) [4], which discovers concepts by clustering segmented image patches, without requiring manual annotation. However, ACE relies on spatial assumptions by considering localized image regions as potential concept candidates, limiting its ability to capture non-local or abstract concepts. A comprehensive survey [18] has reviewed the landscape of concept-based explanations, outlining a range of methodologies and their respective applications.
@@
Another prominent paradigm in XAI is counterfactual explanations, which provide “what-if” scenarios to reveal how minimal semantic changes to an input can alter a model’s prediction [21]. With the emergence of generative models, counterfactual synthesis has advanced significantly, enabling the generation of more realistic and semantically meaningful examples.
@@
In particular, diffusion-based approaches have been proposed to generate counterfactuals by incorporating classifier gradients into the denoising process [22, 23] or by applying adaptive parameterization and cone regularization of gradients [24]. While these methods produce visually realistic counterfactuals, their explanations are typically local to individual samples. In contrast, the Global Counterfactual Directions (GCD) [25] method learns latent directions that consistently invert classifications across samples via a proxy model, enabling more generalizable interpretability.
@@
Recent work has attempted to combine concepts with counterfactual generation. CoLa‑DCE [14] uses classifier guidance to produce spatially constrained, concept‑conditioned counterfactuals. However, these edits remain sample-specific and do not yield reusable global concept vectors. In contrast, Varshney et al. [15] leveraged classifier-guided counterfactual trajectories to identify global disentangled semantic concepts. Despite its strengths, CDCT’s dimension-wise traversal is computationally expensive and limited in detecting complex, multidimensional concepts.
@@
Another line of research leverages clustering in latent spaces to model semantic variation. For example, DifCluE [26] clusters latent embeddings produced by a diffusion autoencoder to generate diverse counterfactual explanations, effectively modeling intra-class variation, but does not explicitly identify global concept directions.
@@
However, most existing approaches reveal several open challenges: reliance on concept annotations or spatial heuristics, a focus on localized or sample-specific edits, the need for exhaustive latent traversal, and limited capacity to capture global, multidimensional semantic concepts. To address these limitations, a novel framework, Concept Directions via Latent Clustering, is proposed as an extension of CDCT [15]. It discovers global, class-specific concept directions by clustering directional latent differences between factual and counterfactual images. This approach enables the extraction of complex, multidimensional concept directions in unsupervised manner and offers a computationally efficient framework for concept-based explanation, enhancing both interpretability and scalability.
@@ 3 Methodology
Figure 1: Overview of the CDLC framework. Counterfactuals are generated using a Latent Diffusion Model (LDM) with classifier guidance, following the procedure used in CDCT [15]. Factual–counterfactual image pairs are encoded using a pretrained Variational Autoencoder (VAE), and the difference between their latent representations is normalized to form unit vectors. These vectors are clustered to identify class-specific concept directions $C_{j}$. During inference, each direction, scaled by a factor $\alpha$, is applied to the test sample’s latent representation to observe its effect on classifier output.
Concept Directions via Latent Clustering (CDLC) is a novel framework designed to extract global, multidimensional concept directions from factual and counterfactual image pairs. It builds upon the counterfactual generation stage of CDCT [15], but instead of relying on computationally expensive dimension-wise latent traversals, it identifies concept directions by clustering unit-normalized difference vectors in a semantic latent space. This section describes the key components of CDLC in detail.
@@ 3.1 Counterfactual Generation via Latent Diffusion Models
Given an input image $x_{f}$ and a trained classifier $f(\cdot)$, a counterfactual image $x_{cf}$ is synthesized to belong to a target class $y_{cf}\neq f(x_{f})$. The counterfactual generation process follows the first stage of the CDCT framework [15], termed Generation of Counterfactual Trajectories, where classifier guidance is integrated into the denoising steps of a latent diffusion model [16] to produce semantically meaningful modifications aligned with the desired class transition. Details of this process can be found in the CDCT framework [15]. Rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. This design choice aims to capture the complete semantic shift required to alter the classifier’s prediction while also reducing storage and computational overhead. The resulting counterfactual is visually realistic and deviates minimally from the factual image, thereby isolating class-specific semantic changes.
@@ 3.2 Discovery of Concept Directions
Each factual image $x_{f}$ and its corresponding counterfactual $x_{cf}$ for a specific target class are encoded using a pretrained Variational Autoencoder, resulting in latent representations $z_{f}=\mathrm{VAE}(x_{f})$ and $z_{cf}=\mathrm{VAE}(x_{cf})$. These latent embeddings capture high-level semantic characteristics of the original and counterfactual images in a compact form.
@@
$\Delta z=z_{cf}-z_{f}$
Each difference vector is reshaped into a flat vector $\Delta z\in\mathbb{R}^{d}$. To focus solely on the directional semantics and eliminate the influence of vector magnitude, each difference vector is normalized to unit length:
$\tilde{z}=\frac{\Delta z}{\|\Delta z\|_{2}}$
@@
A collection of such unit-norm latent difference vectors $\{\tilde{z}_{i}\}_{i=1}^{N}$, derived from multiple factual–counterfactual pairs for a given target class, is clustered into $K$ clusters to identify shared semantic transformations. Cluster centers $C=\{c_{1},c_{2},\ldots,c_{K}\}$ are computed by averaging the unit vectors within each cluster and re-normalizing to unit length:
$c_{k}\in\mathbb{R}^{d},\quad\|c_{k}\|_{2}=1$
Each cluster center $c_{k}$ represents a global concept direction, a consistent and interpretable shift in the latent space associated with the classifier’s decision boundary. These concept directions can then be applied to the latent representations of new test samples to induce semantically meaningful changes in the reconstructed output.
@@ 3.3 Concept Interpretation
To interpret each discovered concept direction $c_{k}$, it is applied to the latent encoding of an unseen test sample:
$z^{\prime}_{\text{test}}=z_{\text{test}}+\alpha\cdot c_{k},$
where $\alpha\in\mathbb{R}$ controls the strength of the semantic manipulation. The modified latent vector $z^{\prime}_{\text{test}}$ is then passed through the VAE decoder to generate a concept-modified image.
In contrast to CDCT, which iteratively modifies each latent dimension to search for influential concepts, CDLC learns global, multi-dimensional concept directions directly from the difference of factual–counterfactual encodings. This leads to a significant increase in computational efficiency and interpretable semantic transformations.
@@ 4 Experiments & Results
This section presents the experimental setup and an analysis of the results produced by Concept Directions via Latent Clustering framework. Both quantitative and qualitative evaluations are performed to assess the semantic coherence of the extracted concept directions and their effectiveness in influencing classifier predictions.
@@ 4.1 Dataset and Classification Model
The same dataset and classification architecture as used in the CDCT framework [15] are adopted for evaluation. Experiments are conducted on a consolidated dermoscopic image dataset derived from the International Skin Imaging Collaboration (ISIC) challenges (2016–2020) [^1] Dataset available at: https://challenge.isic-archive.com/challenges/.. Following the duplicate removal strategy proposed by Cassidy et al. [27], a curated dataset of 29,468 unique images is obtained and stratified into training, validation, and test subsets. The dataset includes eight diagnostic categories: Melanocytic Nevus (NV), Melanoma (MEL), Basal Cell Carcinoma (BCC), Actinic Keratosis (AK), Benign Keratosis (BKL), Dermatofibroma (DF), Vascular Lesions (VASC), and Squamous Cell Carcinoma (SCC).
A ResNet-50 [28] model is trained on the ISIC training partition and used as the target classifier throughout all experiments. This network is used for both generating classifier-guided counterfactuals and evaluating the effectiveness of discovered concept directions.

### v2

@@ Abstract
[…]
@@ 1 Introduction
In high-stakes applications, such as medical diagnosis, financial risk assessment, and autonomous driving, understanding the rationale behind a neural network’s decision is often as important as the decision itself. Explainable Artificial Intelligence (XAI) [arrieta2020explainable, adadi2018peeking] has emerged as a critical research area, aiming to bridge the gap between high-performing black-box models and human interpretability. Among the various XAI paradigms, concept-based explanations [kim2018interpretability, ghorbani2019towards] have gained particular attention due to their ability to express model behavior in terms of high-level, semantically meaningful concepts, rather than low‑level feature weights or pixel-based saliency maps [ribeiro2016should, selvaraju2017grad]. By aligning explanations with concepts recognized by domain experts, these methods facilitate trust [shin2021effects, guo2020explainable], debugging [das2021explainable], and regulatory compliance [ebers2020regulating, panigutti2023role].
@@
Although concept-based explainability has been widely explored using convolutional [kim2018interpretability, ghorbani2019towards] and GAN-based architectures [lang2021explaining, atad2022chexplaining], its application within diffusion-based generative models remains relatively underexplored. Recent works, such as Concept-Guided Latent Diffusion Counterfactual Explanations (CoLa-DCE) [motzkus2024cola], have demonstrated the ability to produce spatially constrained, concept-conditioned counterfactuals for an arbitrary classifier. However, this approach only utilizes concepts to guide counterfactual explanations and does not extract global concept representations that are generalizable across examples.
@@
Varshney et al. [varshney2025generating] proposed the Concept Discovery through Latent Diffusion-based Counterfactual Trajectories (CDCT) framework to discover global concepts that generalize across multiple samples. CDCT generates a classifier-guided counterfactual image trajectory dataset using a latent diffusion model [rombach2022high] and subsequently trains a Variational Autoencoder (VAE) [kingma2013auto] on this trajectory dataset to disentangle classifier-relevant concepts. While CDCT represents a significant advancement in leveraging diffusion-based counterfactuals for concept discovery, it relies on a dimension-wise traversal strategy, wherein each latent variable is modified independently to detect relevant concepts. This exhaustive search procedure is computationally expensive, particularly in high-dimensional latent spaces, and inherently overlooks semantic concepts that arise from interactions among multiple latent dimensions. Consequently, CDCT often identifies only simple concepts, restricting its capacity to uncover high-level semantic directions.
@@
To address these limitations, this paper proposes a novel framework Concept Directions via Latent Clustering (CDLC), that extracts multidimensional concepts by clustering latent difference vectors computed from factual–counterfactual image pairs. These difference vectors, derived from VAE encodings of factual images and their corresponding diffusion-based counterfactuals, capture the classifier-induced transformation in latent space. In contrast to CDCT, which modifies individual latent dimensions, CDLC leverages directional clustering to reveal coordinated latent changes that correspond to semantically meaningful concepts. This approach not only reduces computational complexity by eliminating exhaustive per-dimension search but also enables the discovery of classifier-relevant concept directions that emerge from interactions among multiple latent dimensions, effectively overcoming a limitation of the original CDCT formulation.
@@
    A novel framework, Concept Directions via Latent Clustering (CDLC), is introduced as an extension of CDCT [varshney2025generating], significantly reducing computational complexity compared to dimension-wise search.
@@ 2 Related Work
Explainable Artificial Intelligence (XAI) aims to make machine learning models more transparent by providing human-understandable insights into their decision-making processes [arrieta2020explainable, adadi2018peeking]. Among various XAI approaches, concept-based explanations have gained increasing attention due to their ability to link internal model representations to semantically meaningful, human-interpretable concepts [kim2018interpretability, ghorbani2019towards, poeta2023concept]. In supervised settings, concept-based methods rely on expert-provided annotations, which are costly and time-consuming to collect [kim2018interpretability, lucieri2020explaining]. For instance, Testing with Concept Activation Vectors (TCAV) [kim2018interpretability] requires example sets for each concept and learns linear concept activation vectors to quantify a model’s sensitivity to those concepts. While effective, this reliance on curated concept examples limits the scalability and generalization of such methods. To overcome these limitations, unsupervised concept-based methods have been introduced [ghorbani2019towards, fel2023craft]. A prominent example is Automatic Concept Explanations (ACE) [ghorbani2019towards], which discovers concepts by clustering segmented image patches, without requiring manual annotation. However, ACE relies on spatial assumptions by considering localized image regions as potential concept candidates, limiting its ability to capture non-local or abstract concepts. A comprehensive survey [poeta2023concept] has reviewed the landscape of concept-based explanations, outlining a range of methodologies and their respective applications.
@@
Another prominent paradigm in XAI is counterfactual explanations, which provide “what-if” scenarios to reveal how minimal semantic changes to an input can alter a model’s prediction [wachter2017counterfactual]. With the emergence of generative models, counterfactual synthesis has advanced significantly, generating more realistic and semantically meaningful examples.
@@
In particular, diffusion-based approaches have been proposed to generate counterfactuals by incorporating classifier gradients into the denoising process [jeanneret2022diffusion, jeanneret2023adversarial] or by applying adaptive parameterization and cone regularization of gradients [augustin2022diffusion]. While these methods produce visually realistic counterfactuals, their explanations are typically local to individual samples. In contrast, the Global Counterfactual Directions (GCD) [sobieski2024global] method learns latent directions that consistently invert sample classifications via a proxy model, enabling more generalizable interpretability.
@@
Recent work has attempted to combine concepts with counterfactual generation. CoLa‑DCE [motzkus2024cola] uses classifier guidance to produce spatially constrained, concept‑conditioned counterfactuals. Similarly, DiffEx [kazimi2025explaining] combines a vision–language model with diffusion semantics to generate a semantic hierarchy of attributes and rank their influence on classifier decisions. However, these methods remain sample-specific and do not yield reusable global concept vectors. In contrast, Varshney et al. [varshney2025generating] leveraged classifier-guided counterfactual trajectories to identify global disentangled semantic concepts. Despite its strengths, CDCT’s dimension-wise traversal is computationally expensive and limited in detecting complex, multidimensional concepts.
@@
Another line of research leverages clustering in latent spaces to model semantic variation. For example, DifCluE [jain2025difclue] clusters latent embeddings produced by a diffusion autoencoder to generate diverse counterfactual explanations, effectively modeling intra-class variation, but does not explicitly identify global concept directions. A parallel line clusters concept-attribution vectors rather than latent features. PCX [dreyer2024understanding] computes per-sample concept-relevance vectors and fits class-wise Gaussian mixture models to cluster them into prototypical decision strategies that explain model behavior.
@@
However, most existing approaches reveal several open challenges: reliance on concept annotations or spatial heuristics, a focus on localized or sample-specific edits, the need for exhaustive latent traversal, and limited capacity to capture global, multidimensional semantic concepts. To address these limitations, a novel framework, Concept Directions via Latent Clustering, is proposed as an extension of CDCT [varshney2025generating]. It discovers global, class-specific concept directions by clustering directional latent differences between factual and counterfactual images. This approach enables the extraction of complex, multidimensional concept directions in an unsupervised manner and offers a computationally efficient framework for concept-based explanation, enhancing both interpretability and scalability.
@@ 3 Methodology
Figure 1: Overview of the CDLC framework. Counterfactuals are generated using a Latent Diffusion Model (LDM) with classifier guidance, following the procedure used in CDCT [varshney2025generating]. Factual–counterfactual image pairs are encoded using a pretrained Variational Autoencoder (VAE), and the difference between their latent representations is normalized to form unit vectors. These vectors are clustered to identify class-specific concept directions $C_{j}$. During inference, each direction, scaled by a factor $\alpha$, is applied to the test sample’s latent representation to observe its effect on classifier output.
Concept Directions via Latent Clustering (CDLC) is a novel framework designed to extract global, multidimensional concept directions from factual and counterfactual image pairs. It builds upon the counterfactual generation stage of CDCT [varshney2025generating]. However, instead of relying on computationally expensive dimension-wise latent traversals, it identifies concept directions by clustering unit-normalized difference vectors in a semantic latent space. This section describes the key components of CDLC in detail.
@@ 3.1 Counterfactual Generation via Latent Diffusion Models
Given an input image $x_{f}$ and a trained classifier $f(\cdot)$, a counterfactual image $x_{cf}$ is synthesized to belong to a target class $y_{cf}\neq f(x_{f})$. The counterfactual generation process follows the first stage of the CDCT framework [varshney2025generating], termed Generation of Counterfactual Trajectories, where classifier guidance is integrated into the denoising steps of a latent diffusion model [rombach2022high] to produce semantically meaningful modifications aligned with the desired class transition. Details of this process can be found in the CDCT framework [varshney2025generating]. Rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. This design choice aims to capture the semantic shift required to alter the classifier’s prediction while reducing storage and computational overhead. The resulting counterfactual is visually realistic and deviates minimally from the factual image, thereby isolating class-specific semantic changes.
@@ 3.2 Discovery of Concept Directions
Each factual image $x_{f}$ and its corresponding counterfactual $x_{cf}$ for a specific target class are encoded using a pretrained Variational Autoencoder, resulting in latent representations $\mathbf{z}_{f}=\mathrm{VAE}(x_{f})$ and $\mathbf{z}_{cf}=\mathrm{VAE}(x_{cf})$. These latent embeddings capture high-level semantic characteristics of the original and counterfactual images in a compact form.
@@
$\Delta\mathbf{z}=\mathbf{z}_{cf}-\mathbf{z}_{f}$
Each difference vector is reshaped into a flat vector $\Delta\mathbf{z}\in\mathbb{R}^{d}$. To focus solely on the directional semantics and eliminate the influence of vector magnitude, each difference vector is normalized to unit length:
$\tilde{\mathbf{z}}=\frac{\Delta\mathbf{z}}{\|\Delta\mathbf{z}\|_{2}}$
@@
A collection of such unit-norm latent difference vectors $\{\tilde{\mathbf{z}}_{i}\}_{i=1}^{N}$, derived from multiple factual–counterfactual pairs for a given target class, is clustered into $K$ clusters to identify shared semantic transformations. Cluster centers $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$ are computed by averaging the unit vectors within each cluster and re-normalizing to unit length:
$\mathbf{c}_{k}\in\mathbb{R}^{d},\quad\|\mathbf{c}_{k}\|_{2}=1$
Each cluster center $\mathbf{c}_{k}$ represents a global concept direction, a consistent and interpretable shift in the latent space associated with the classifier’s decision boundary. These concept directions can then be applied to the latent representations of new test samples to induce semantically meaningful changes in the reconstructed output.
@@ 3.3 Concept Interpretation
To interpret each discovered concept direction $\mathbf{c}_{k}$, it is applied to the latent encoding of an unseen test sample:
$\mathbf{z}^{\prime}_{\text{test}}=\mathbf{z}_{\text{test}}+\alpha\cdot\mathbf{c}_{k},$
where $\alpha\in\mathbb{R}$ controls the strength of the semantic manipulation. The modified latent vector $\mathbf{z}^{\prime}_{\text{test}}$ is then passed through the VAE decoder to generate a concept-modified image.
In contrast to CDCT, which iteratively modifies each latent dimension to search for influential concepts, CDLC learns global, multidimensional concept directions directly from the difference of factual–counterfactual encodings. CDLC avoids storing intermediate image trajectories and relies on CPU-only clustering of latent differences; no auxiliary VAE training or iterative dimension-wise search is required. These changes lead to a significant increase in computational efficiency and interpretable semantic transformations.
@@ 4 Experiments & Results
This section presents the experimental setup and an analysis of the results produced by Concept Directions via Latent Clustering framework. Quantitative and qualitative evaluations assess the semantic coherence of the extracted concept directions and their effectiveness in influencing classifier predictions.
@@ 4.1 Dataset and Classification Model
The same dataset and classification architecture used in the CDCT framework [varshney2025generating] are adopted for evaluation. Experiments are conducted on a consolidated dermoscopic image dataset derived from the International Skin Imaging Collaboration (ISIC) challenges (2016–2020) [^1] Dataset available at: https://challenge.isic-archive.com/challenges/.. Following the duplicate removal strategy proposed by Cassidy et al. [cassidy2022analysis], a curated dataset of 29,468 unique images is obtained and stratified into training, validation, and test subsets. The dataset includes eight diagnostic categories: Melanocytic Nevus (NV), Melanoma (MEL), Basal Cell Carcinoma (BCC), Actinic Keratosis (AK), Benign Keratosis (BKL), Dermatofibroma (DF), Vascular Lesions (VASC), and Squamous Cell Carcinoma (SCC).
A ResNet-50 [he2016deep] model is trained on the ISIC training partition and used as the target classifier throughout all experiments. This network is used to generate classifier-guided counterfactuals and evaluate the effectiveness of discovered concept directions.